OpenStudy (anonymous):

If x+28/7=8 then what is the value of x?

OpenStudy (anonymous):

What will you get if you divide 28 by 7??

OpenStudy (callisto):

Is the question (i) \(\frac{x+28}{7}=8\) or (ii) \(x+\frac{28}{7}=8\) ?

OpenStudy (anonymous):

@waterineyes 4

OpenStudy (anonymous):

Firstly answer to callisto question..

OpenStudy (anonymous):

@Callisto (i)

OpenStudy (anonymous):

Multiply both sides by 7 and tell me what you get..

OpenStudy (anonymous):

@waterineyes i do not get it

OpenStudy (anonymous):

See if you multiply both sides by 7 like this: \[7 (\frac{x + 28}{7}) = 7 \times 8\] Now solve..

OpenStudy (anonymous):

56?

OpenStudy (anonymous):

This is on right hand side.. What you left with left hand side?? Hint: 7 and 7 can get cancel..

OpenStudy (anonymous):

nothing

OpenStudy (anonymous):

i do not have anything on the left hand side

OpenStudy (anonymous):

\[7(\frac{x + 28}{7})\] This is your left hand side..

OpenStudy (anonymous):

yes but i cancel out the 7's

OpenStudy (anonymous):

So you left with the numerator.. Isn't it??

OpenStudy (anonymous):

ok, i have 28!

OpenStudy (anonymous):

x+28 is all that is left

OpenStudy (anonymous):

Only 28??

OpenStudy (anonymous):

Yes now you are right.. So all you with this: \[x+28 = 56\]

OpenStudy (anonymous):

Now Subtract 28 from both the sides to find x. Do it..

OpenStudy (anonymous):

that gives me 28

OpenStudy (anonymous):

which is the answer, omg thank you so much

OpenStudy (anonymous):

Welcome dear...